# columns=colsep=4pt, rows=rowsep=.5pt

### v1

@@ Abstract
Prototypical self-supervised learning methods consistently suffer from partial prototype collapse, where multiple prototypes converge to nearly identical representations. This undermines their central purpose—providing diverse and informative targets to guide encoders toward rich representations—and has led practitioners to over-parameterize prototype sets or add ad-hoc regularizers, which mitigate symptoms rather than address the root cause. We empirically trace the collapse to the joint optimization of encoders and prototypes, which encourages a type of shortcut learning: early in training prototypes drift toward redundant representations that minimize loss without necessarily enhancing representation diversity. To break the joint optimization, we introduce a fully decoupled training strategy that learns prototypes and encoders under separate objectives. Concretely, we model prototypes as a Gaussian mixture updated with an online EM-style procedure, independent of the encoder’s loss. This simple yet principled decoupling eliminates prototype collapse without explicit regularization and yields consistently diverse prototypes and stronger downstream performance.
@@ 1 Introduction
Prototypical self-supervised learning (SSL) (Siméoni et al., 2025; Caron et al., 2021) have come to rival the effectiveness of language-supervised alternatives in representation learning (Fan et al., 2025). Yet recent work has demonstrated that many prototypical SSL approaches suffer from a phenomenon known as partial prototype collapse, in which multiple prototypes converge to indistinguishable representations (Govindarajan et al., 2023, 2024). This phenomenon could explain why over-parameterizing the number of prototypes used has been a popular choice to improve performance in recent prototypical frameworks (Venkataramanan et al., 2025; Oquab et al., 2024; Siméoni et al., 2025).
@@
Our experiments show that the practice of jointly optimizing the encoder and prototypes contributes to this collapse. The intuition is that joint optimization drives prototypes towards redundant representations early in training by exploiting shortcuts that minimize the loss at the cost of diverse and semantically meaningful representations. This leads to prototype over-parametrization; an expensive solution that mitigates symptoms without directly addressing the underlying issues (Venkataramanan et al., 2025; Oquab et al., 2024; Siméoni et al., 2025) while impeding a model’s ability to learn stronger representations and diminishing robustness to imbalanced data distributions (Govindarajan et al., 2024; Wen et al., 2024).
@@
We find that these effects are particularly pronounced in the instance-level formulations of prototypical SSL. In Fig. LABEL:fig:unif-dino, we reaffirm previous findings (Govindarajan et al., 2024), showing how DINO’s (Caron et al., 2021) final prototype distribution tends to collapse to one or two modes, which severely limits the diversity of representations. This phenomenon also persists in more recent methods such as DINOv2 (Oquab et al., 2024), cf. Fig. LABEL:fig:unif-dinov2. While recent efforts have looked to address this issue, either via explicit regularization (Govindarajan et al., 2024), cf. Fig. LABEL:fig:unif-dino-kp, or implicit mechanisms (Silva and Ramírez Rivera, 2023), cf. Fig. LABEL:fig:unif-carp, these models still exhibit collapse, suggesting that the underlying issue is not yet fully resolved.
Motivated by these observations, we introduce a decoupling strategy designed to break the joint-optimization procedure. In our approach, the encoder and prototypes are learned under separate objectives—see Fig. 2. Concretely, we model the prototypes as a Gaussian mixture, updated via an expectation maximization (EM) independently of the encoder’s loss. This decoupled training eliminates partial prototype collapse, cf. Fig. LABEL:fig:unif-carp-dec, and points to joint optimization as the most likely culprit driving partial prototype collapse.
@@
Our main contributions are as follows: (i) we conduct a systematic and quantitative analysis of a broad range of prototypical SSL frameworks, demonstrating that partial prototype collapse extends well beyond the DINO family of models; (ii) we identify the underlying mechanism driving this collapse as the joint optimization of encoders and prototypes under a shared loss, which encourages redundant prototype representations; and (iii) we introduce a fully decoupled training framework that isolates prototype estimation from encoder learning, achieving consistently high prototype diversity throughout training and yielding stronger representations with improved robustness to long-tailed data distributions.
@@ 2.1 Preliminaries
Prototypical Self-Supervised Formulation. Among recent SSL approaches, prototypical formulations (where network outputs are aligned to a set of learnable, class-agnostic prototypes) have achieved state-of-the-art performance on several vision benchmarks (Siméoni et al., 2025). We begin by describing the general prototypical framework used throughout this paper. Our exposition emphasizes the instance-level objective (Caron et al., 2021; Silva and Ramírez Rivera, 2022, 2023), but the same formulation can readily be extended to incorporate a dense objective (Zhou et al., 2022) or reformulated with an explicit dense prediction objective (Darcet et al., 2025).
Given $J$ stochastic augmentations $\{v^{j}\}_{j=1}^{J}$ of an image $x$, a student backbone with projection head $f_{\theta}$ and an EMA-updated teacher $f_{\phi}$ map each view to a latent representation $h^{j}_{(\cdot)}=f_{(\cdot)}(v^{j})\in\mathbb{R}^{D}$. Similarity scores with a learnable prototype set $C=[c_{1},c_{2},\dots,c_{K}]\in\mathbb{R}^{D\times K}$ are computed as $z_{(\cdot)}^{j}=h_{(\cdot)}^{j}C^{\top}\in\mathbb{R}^{K}$ and converted into prototype-assignment probabilities via a softmax with an optional temperature parameter $\tau>0$ through
@@
Partial Prototype Collapse. Govindarajan et al. (2024) investigated the DINO-family of methods (Caron et al., 2021; Zhou et al., 2022; Assran et al., 2022, 2023a) and demonstrated that these prototypical SSL approaches are susceptible to partial prototype collapse. This phenomenon arises when multiple prototypes converge to nearly identical representations during training. While this behavior does not amount to a complete collapse of all prototypes, it, nevertheless, reduces the diversity of representations and may hinder the effectiveness of downstream tasks. To formalize this behavior, the authors introduced the following definition of partial collapse.
@@
Applying this definition, they further showed that the DINO-family of methods retained only about 2%–40% of their initially initialized prototypes as unique, thereby, revealing a substantial redundancy in the learned representations. To alleviate this issue, they proposed applying KoLeo regularization (Beirlant et al., 1997; Delattre and Fournier, 2017; Sablayrolles et al., 2019) directly to the prototypes, effectively introducing a diversity-enforcing auxiliary term into the objective. While this helped reduce collapse, it did not fully solve the problem. Their approach also introduces a new hyperparameter, creating a delicate trade-off: insufficient regularization limits diversity, whereas excessive regularization impair encoder learning. While the work of Govindarajan et al. (2024) represented an important first step in identifying partial prototype collapse, their analysis did not examine its underlying mechanisms and was limited to methods within the DINO family. This opens the door to a broader line of inquiry: Is partial prototype collapse restricted to the DINO family of methods, and what mechanisms drive this type of collapse?
@@ 3 Solving Partial Prototype Collapse through Decoupling
If partial prototype collapse truly stems from the joint optimization of prototypes and encoders, then separating their updates should prevent the collapse. Motivated by this intuition, and unlike CAPI (Darcet et al., 2025), which decouples only the teacher’s prototypes from the main loss while keeping the student prototypes jointly optimized, we propose a fully decoupled training procedure that isolates prototype estimation from encoder learning, cf. Fig. LABEL:fig:dec-jea, allowing us to directly probe the role of joint optimization in driving collapse.
@@
$C^{t}=\operatorname*{arg\,min}_{C\in\mathcal{C}}\;\mathcal{L}_{C}\left(C^{t-1};h^{t}_{\phi}\right),$ (3)
@@
$\theta^{t+1}=\operatorname*{arg\,min}_{\theta}\;\mathcal{L}_{f}\left(h^{t},C^{t}\right),$ (4)
where $h^{t}$ denotes the latent representations at iteration $t$, and $\mathcal{L}_{C}$ is a loss that estimates prototypes directly from the latent features, independent of the encoder’s loss $\mathcal{L}_{f}$. By fully separating prototype estimation from encoder optimization, our method removes the shortcut incentive entirely and aims to achieve stable, diverse prototypes throughout training.
@@ 4.1 Experimental Setup
To evaluate our decoupling strategy, we adopt CARP (Silva and Ramírez Rivera, 2023). We select CARP for two main reasons. First, it is an instance-based approach, which we have shown to exhibit a higher degree of prototypical collapse (cf. Table 1) which offers a more challenging setting to evaluate the efficacy of our decoupling method, as the collapse is more pronounced, while also isolating the effect of MIM objectives on prototype diversity, which we leave for future work. Second, extensive ablation experiments have demonstrated that CARP maintains robust and stable performance across a wide range of hyperparameters (Silva and Ramírez Rivera, 2023), making it a suitable framework for evaluating our decoupling strategy without extensive hyperparameter tuning. [^5] This contrasts with the DINO formulation, which has been shown in several studies to suffer from substantial hyperparameter sensitivity (Ruan et al., 2023; Wu et al., 2025).
@@ 4.3 Training dynamics
Although previous studies have examined partial prototype collapse, they have focused almost exclusively on its manifestation at the end of training (Govindarajan et al., 2023, 2024), leaving its interaction with the training dynamics largely unexplored. To address this and gain deeper insight into the relationship between prototype collapse and training progression, in Fig. 4, we track the number of unique prototypes in use and the linear evaluation accuracy on ImageNet-1k (Deng et al., 2009) over the first 100 epochs for multiple methods. Specifically, we include CARL (Silva and Ramírez Rivera, 2022) and CARP (Silva and Ramírez Rivera, 2023) (CARP being an extension of the CARL framework that introduces random partitioning) alongside DINO (Caron et al., 2021), DINO + KoLeoProto (Govindarajan et al., 2024), and our proposed decoupled approach.
A first observation is that partial prototype collapse emerges very early in training: after only 10 epochs, two-thirds of the prototypes have already collapsed. Introducing the KoLeo-Proto (KP) regularizer (Govindarajan et al., 2024) proves effective at maintaining prototype diversity, underscoring its role as a strong collapse-prevention mechanism. While DINO+KP exhibits a modest drop in accuracy during the early stages of training, this effect diminishes over time, and the final model slightly outperforms vanilla DINO. This suggests that preserving higher prototype diversity throughout training ultimately leads to stronger representations in the converged model.
@@
Finally, we find that CARP + Decoupling maintains full prototype utilization while outperforming all other methods in the early and late phases of training, indicating that decoupling does not hinder early training dynamics. By the end of training, CARP + Decoupling improves upon CARP, demonstrating that enhanced prototype diversity yields stronger representations.
@@ 4.4 Effect of prototype diversity across head, medium and tail classes
Methods
Head
(>100>100)
Medium
(>20&100>20\ $\leq$ 100)
Tail
($\leq 20$)
All
@@ Methods
CARP + Decoupling 59.1 $\uparrow$  3.1 49.3 $\uparrow$  2.4 45.9 $\uparrow$  3.4 48.9 $\uparrow$  3.0
@@
Table 2 shows that greater prototype diversity substantially enhances robustness in long-tailed datasets. While DINO exhibits a 3.7-percentage-point improvement attributable to the increase in prototype diversity achieved through KP regularization, our proposed decoupling approach on top of CARP yields an improvement of $3.0$ percentage points, although CARP already enjoys greater robustness on long-tailed data due to its higher prototype diversity than DINO. Furthermore, we find that all examined distributions of class frequencies benefit from increased prototype diversity. This finding underscores the necessity of mitigating prototypical collapse in prototypical SSL frameworks and suggests that integrating such mechanisms may enhance the robustness of methods trained on less curated data. Finally, CARP + Decoupling surpasses DINO + KP in tail-class accuracy by nearly one percentage point, this combined with the substantial gains obtained through decoupling the joint optimization in CARP, highlights the efficacy of our decoupling strategy in enhancing prototype diversity.

### v2

# Why Prototypes Collapse: Diagnosing and Preventing Partial Collapse in Prototypical Self-Supervised Learning

@@ Abstract
Prototypical self-supervised learning methods consistently suffer from partial prototype collapse, where multiple prototypes converge to nearly identical representations. This undermines their central purpose—providing diverse and informative targets to guide encoders toward rich representations—and has led practitioners to over-parameterize prototype sets or add ad-hoc regularizers, which mitigate symptoms rather than address the root cause. We empirically trace the collapse to the joint optimization of encoders and prototypes, which encourages a type of shortcut learning: early in training prototypes drift toward redundant representations that minimize loss without necessarily enhancing representation diversity. To break the joint optimization, we introduce a fully decoupled training strategy that learns prototypes and encoders under separate objectives. Concretely, we model prototypes as a Gaussian mixture updated with an online EM-style procedure, independent of the encoder’s loss. This simple yet principled decoupling eliminates prototype collapse without explicit regularization and yields consistently diverse prototypes, which in several settings translate to improved downstream performance.
@@ 1 Introduction
Prototypical self-supervised learning (SSL) frameworks (Siméoni et al., 2025; Caron et al., 2021) in the image domain have come to rival the effectiveness of language-supervised alternatives in representation learning (Fan et al., 2025). In these frameworks, prototypes are learnable vectors that act as cluster anchors, guiding learning so that samples organize into semantically coherent regions of the embedding space. However, recent work has demonstrated that many prototypical SSL approaches suffer from a phenomenon known as partial prototype collapse (formal definition in Def. 2.1), in which multiple prototypes converge to indistinguishable representations (Govindarajan et al., 2023; 2024). This phenomenon could explain why over-parameterizing the number of prototypes used has been a popular choice to improve performance in recent prototypical frameworks (Venkataramanan et al., 2025; Oquab et al., 2024; Siméoni et al., 2025).
@@
Our experiments show that the practice of jointly optimizing the encoder and prototypes contributes to this collapse. The intuition is that joint optimization drives prototypes towards redundant representations early in training by exploiting shortcuts that minimize the loss at the cost of diverse and semantically meaningful representations. This encourages the use of larger prototype sets; an expensive solution that mitigates symptoms without directly addressing the underlying issues (Venkataramanan et al., 2025; Oquab et al., 2024; Siméoni et al., 2025) while impeding a model’s ability to learn stronger representations and diminishing robustness to imbalanced data distributions (Govindarajan et al., 2024; Wen et al., 2024).
@@
We find that these effects are particularly pronounced in the instance-level formulations of prototypical SSL. In Fig. LABEL:fig:unif-dino, we reaffirm previous findings (Govindarajan et al., 2024), showing how DINO’s (Caron et al., 2021) final prototype distribution tends to collapse to one or two modes, which severely limits the diversity of representations. This phenomenon also persists in more recent methods such as DINOv2 (Oquab et al., 2024), cf. Fig. LABEL:fig:unif-dinov2. While recent efforts have looked to address this issue, either via explicit regularization (Govindarajan et al., 2024), cf. Fig. LABEL:fig:unif-dino-kp, or implicit mechanisms (Silva and Ramírez Rivera, 2023), cf. Fig. LABEL:fig:unif-carp, these models still exhibit collapse, suggesting that the underlying issue is not yet fully resolved. Motivated by these observations, we introduce a decoupling strategy designed to break the joint-optimization procedure. In our approach, the encoder and prototypes are learned under separate objectives—see Fig. 2. Concretely, we model the prototypes as a Gaussian mixture, updated via an expectation maximization (EM) independently of the encoder’s loss. This decoupled training eliminates partial prototype collapse, cf. Fig. LABEL:fig:unif-carp-dec, and points to joint optimization as the most likely culprit driving partial prototype collapse.
@@
Our main contributions are as follows: (i) we conduct a systematic and quantitative analysis of a broad range of prototypical SSL frameworks, demonstrating that partial prototype collapse extends well beyond the DINO family of models; (ii) we identify the underlying mechanism driving this collapse as the joint optimization of encoders and prototypes under a shared loss, which encourages redundant prototype representations; and (iii) we introduce a fully decoupled training framework that isolates prototype estimation from encoder learning, achieving consistently high prototype diversity throughout training, with downstream representation quality and robustness that improve in several prototypical SSL settings.
@@ 2.1 Preliminaries
Prototypical Self-Supervised Formulation. Among recent SSL approaches, prototypical formulations (where network outputs are aligned to a set of learnable, class-agnostic prototypes) have achieved state-of-the-art performance on several vision benchmarks (Siméoni et al., 2025). We begin by describing the general prototypical framework used throughout this paper. Our exposition emphasizes the instance-level objective (Caron et al., 2021; Silva and Ramírez Rivera, 2022; 2023), but the same formulation can readily be extended to incorporate a dense objective (Zhou et al., 2022) or reformulated with an explicit dense prediction objective (Darcet et al., 2025).
Given $J$ stochastic augmentations $\{v^{j}\}_{j=1}^{J}$ of an image $x$, a student backbone with projection head $f_{\theta}$ and an Exponential Moving Average (EMA)-updated teacher $f_{\phi}$ map each view to a latent representation $h^{j}_{(\cdot)}=f_{(\cdot)}(v^{j})\in\mathbb{R}^{D}$. Similarity scores with a learnable prototype set $C\in\mathcal{C}:=\mathbb{R}^{D\times K}$ are computed as $z_{(\cdot)}^{j}=h_{(\cdot)}^{j}C\in\mathbb{R}^{K}$ and converted into prototype-assignment probabilities via a softmax with an optional temperature parameter $\tau>0$ through
@@
Partial Prototype Collapse. Govindarajan et al. (2024) investigated the DINO-family of methods (Caron et al., 2021; Zhou et al., 2022; Assran et al., 2022; 2023b) and demonstrated that these prototypical SSL approaches are susceptible to partial prototype collapse. This phenomenon arises when multiple prototypes converge to nearly identical representations during training. While this behavior does not amount to a complete collapse of all prototypes, it, nevertheless, reduces the diversity of representations and may hinder the effectiveness of downstream tasks. To formalize this behavior, the authors introduced the following definition of partial collapse.
@@
Applying this definition, they further showed that the DINO-family of methods retained only about 2%–40% of their initially initialized prototypes as unique, thereby, revealing a substantial redundancy in the learned representations. To alleviate this issue, they proposed applying KoLeo regularization (Beirlant et al., 1997; Delattre and Fournier, 2017; Sablayrolles et al., 2019) directly to the prototypes. Specifically, they added a KoLeo-Prototype (KP) loss term, $\mathcal{L}_{\text{KP}}(C)=-\frac{1}{K}\sum_{k=1}^{K}\log(d_{k})$, where $d_{k}=\min_{i\neq k}\lVert c_{k}-c_{i}\rVert$, as an auxiliary term to encourage prototype diversity. While this helped reduce collapse, it did not fully solve the problem. Their approach also introduced a new hyperparameter $\lambda_{\text{KP}}$, which balances the KP loss with the overall objective, creating a delicate trade-off: insufficient regularization limits diversity, whereas excessive regularization can impair encoder learning.
While the work of Govindarajan et al. (2024) represented an important first step in identifying partial prototype collapse, their analysis did not examine its underlying mechanisms and was limited to methods within the DINO family. This opens the door to a broader line of inquiry: Is partial prototype collapse restricted to the DINO family of methods, and what mechanisms drive this type of collapse?
@@ 3 Solving Partial Prototype Collapse through Decoupling
If partial prototype collapse is influenced by the joint optimization of prototypes and encoders, then separating their updates may reduce the likelihood or severity of the collapse. Motivated by this intuition, and unlike CAPI (Darcet et al., 2025), which decouples only the teacher’s prototypes from the main loss while keeping the student prototypes jointly optimized, we propose a fully decoupled training procedure that isolates prototype estimation from encoder learning, cf. Fig. LABEL:fig:dec-jea, allowing us to directly probe the role of joint optimization in driving collapse.
@@
$C^{t}=\operatorname*{arg\,min}_{C\in\mathcal{C}}\;\mathcal{L}_{C}\left(C^{t-1},h^{t}_{\phi}\right),$ (3)
@@
$\theta^{t+1}=\operatorname*{arg\,min}_{\theta}\;\mathcal{L}_{f}\left(h^{t}_{\theta},C^{t}\right),$ (4)
where $h^{t}$ denotes the latent representations at iteration $t$, and $\mathcal{L}_{C}$ is a loss that estimates prototypes directly from the latent features, independent of the encoder’s loss $\mathcal{L}_{f}$. By fully separating prototype estimation from encoder optimization, our method removes the shortcut incentive and yields stable, diverse prototypes throughout training. A theoretical motivation is provided in Appendix C.1.
@@ 4.1 Experimental Setup
To evaluate our decoupling strategy, we adopt CARP (Silva and Ramírez Rivera, 2023) as our primary point of reference. We select CARP for two main reasons. First, because it is an instance-based approach, and instance-based methods in general exhibit a higher degree of prototypical collapse (cf. Table 1), CARP provides a more challenging setting for evaluating our decoupling method, while also isolating the effect of MIM objectives on prototype diversity, which we leave for future work. Second, extensive ablation experiments have demonstrated that CARP maintains robust and stable performance across a wide range of hyperparameters (Silva and Ramírez Rivera, 2023), making it a suitable framework for evaluating our decoupling strategy without extensive hyperparameter tuning. In addition, we evaluate DINO (Caron et al., 2021) with our decoupling method, incorporating DINO enables a broader assessment of the generality of our method and its impact on prototypical diversity.
@@ 4.3 Training dynamics
Although previous studies have examined partial prototype collapse, they have focused almost exclusively on its manifestation at the end of training (Govindarajan et al., 2023; 2024), leaving its interaction with the training dynamics largely unexplored. To address this and gain deeper insight into the relationship between prototype collapse and training progression, in Fig. 4, we track the number of unique prototypes in use and the linear evaluation accuracy on ImageNet-1k (Deng et al., 2009) over the first 100 epochs for multiple methods. Specifically, we include CARL (Silva and Ramírez Rivera, 2022) and CARP (Silva and Ramírez Rivera, 2023) (CARP being an extension of the CARL framework that introduces random partitioning) alongside DINO (Caron et al., 2021), DINO + KoLeoProto (KP) (Govindarajan et al., 2024), and our proposed decoupled approach.
A first observation is that partial prototype collapse emerges very early in training: after only 10 epochs, two-thirds of the prototypes have already collapsed. Introducing the KP regularizer (Govindarajan et al., 2024) proves effective at maintaining prototype diversity, underscoring its role as a strong collapse-prevention mechanism. While DINO+KP exhibits a modest drop in accuracy during the early stages of training, this effect diminishes over time, and the final model slightly outperforms vanilla DINO. This suggests that preserving higher prototype diversity throughout training ultimately leads to stronger representations in the converged model.
@@
Finally, we find that CARP + Decoupling maintains full prototype utilization while outperforming all other methods in the early and late phases of training, indicating that decoupling does not hinder its early training dynamics. By the end of training, CARP + Decoupling improves upon CARP, demonstrating that enhanced prototype diversity yields stronger representations.
@@ 4.4 Effect of prototype diversity across head, medium and tail classes
Methods Head (>100>100) Medium (>20&100>20\ $\leq$ 100) Tail ($\leq 20$) All
@@
DINO + Decoupling 44.9 $\downarrow$  10.3 37.3 $\downarrow$  8.9 33.1 $\downarrow$  8.8 36.2 $\downarrow$  9.1
@@
CARP + KP 55.5 $\downarrow$  0.5 46.4 $\downarrow$  0.5 41.8 $\downarrow$  0.7 45.3 $\downarrow$  0.6
[…]
@@
Table 2 presents the performance of all methods across the head, medium, and tail splits of iNaturalist 2018. Consistent with prior observations (Govindarajan et al., 2024), increasing prototype diversity often improves robustness under long-tailed class distributions. In addition, this experiment shows that these gains are not confined to head classes: improvements appear across all class-frequency regimes, including the tail classes. For DINO, KP regularization yields a 3.7-percentage-point gain in overall accuracy. In contrast, applying our decoupling strategy directly to DINO leads to a substantial performance drop across all class-frequency regimes. We attribute this decline to the high sensitivity of DINO to excessive prototype diversity; in ImageNet-1k, extensive hyperparameter tuning was required to stabilize DINO under decoupling, and these settings do not transfer effectively to the heavily imbalanced iNaturalist domain. A detailed analysis of DINO’s hyperparameter sensitivity is provided in Appendix C.3.
For CARP, the trends differ: adding KP results in a slight decrease in accuracy, which we suspect is linked to the choice of KP regularization strength. Our experiments adopt the value recommended by Govindarajan et al. (2024), selected based on ablations conducted for DINO, an architecture exhibiting substantially stronger prototypical collapse than CARP, as demonstrated in Figure LABEL:fig:unique-protos-v-eps. When applied to CARP under long-tailed conditions, using the same strength appears to hinder learning. In contrast, integrating our decoupling strategy with CARP yields consistent improvements across all frequency regimes, amounting to a 3.0-percentage-point gain overall.
In summary, and in line with Govindarajan et al. (2024), increasing prototype diversity in prototypical SSL frameworks may substantially improve robustness under long-tailed data distributions. Our granular analysis confirms that these gains extend consistently to tail classes rather than being driven solely by improved modeling of head classes. However, for DINO, an excessively large degree of prototype diversity appears to render optimization unstable and requires extensive hyperparameter adjustment. Likewise, for KP, our experiments indicate that careful tuning of the regularization strength across frameworks is necessary when operating in long-tailed settings.